# C-3DPO: Constrained Controlled Classification for Direct Preference Optimization

### v1

@@ 5 Constrained Controlled Classification for DPO (C-3DPO)
We now present a general family of constraints to fix the under-specified nature of the DPO-style algorithms described in Section 4. These constraints are on the probability mass of the winner-loser pair and we use them to control how much this mass changes from the reference policy, $\pi_{\textrm{ref}}$, to target policy $\pi_{\theta}$. We then show how these constraints can be incorporated into any DPO-style loss function and propose our algorithms which we refer to as Constrained Controlled Classification for Direct Preference Optimization (C-3DPO).
@@
The constraint, with respect to an arbitrary function $\varphi:\mathbb{R}\rightarrow\mathbb{R}$, takes the general form:
$\varphi\big{(}\pi_{\theta}(y_{w}|x)\big{)}$ $+\varphi\big{(}\pi_{\theta}(y_{l}|x)\big{)}$ (10)
$=\varphi\big{(}\pi_{\textrm{ref}}(y_{w}|x)\big{)}+\varphi\big{(}\pi_{\textrm{ref}}(y_{l}|x)\big{)}\ .$
Note that the right-hand-side is fixed during training, so the two probabilities on the left-hand-side cannot move in the same direction.
@@
We now generalize this intuition by showing that when the constraint function $\varphi$ is monotonic and is added to the solution characterization of DPO-style algorithms (9), then we can control the direction of the movement of probability mass for winner-loser pairs.
@@
Proof.
For simplicity we assume that $\varphi$ is monotonically increasing (the same arguments can be easily applied to the monotonically decreasing case). From (9), we may write
$\frac{\pi_{\theta^{*}}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{w}|x)}=\sqrt[\beta]{\frac{1-\varepsilon}{\varepsilon}}\cdot\frac{\pi_{\theta^{*}}(y_{l}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}\ .$
Since $\varepsilon<1/2$, we have that $(1-\varepsilon)/\varepsilon>1$, and therefore
We will show that $\pi_{\theta^{*}}(y_{l}|x)<\pi_{\textrm{ref}}(y_{l}|x)$ by contradiction. To this end, we assume that $\pi_{\theta^{*}}(y_{l}|x)\geq\pi_{\textrm{ref}}(y_{l}|x)$. From (11), we immediately obtain that $\pi_{\theta^{*}}(y_{w}|x)>\pi_{\textrm{ref}}(y_{w}|x)$. Applying the constraint $\varphi$ to the following two inequalities:
and using its monotonicity, we obtain
$\varphi\big{(}\pi_{\theta^{*}}(y_{w}|x)\big{)}$ $>$ $\varphi\big{(}\pi_{\textrm{ref}}(y_{w}|x)\big{)}$
$\varphi\big{(}\pi_{\theta^{*}}(y_{l}|x)\big{)}$$\geq$ $\varphi\big{(}\pi_{\textrm{ref}}(y_{l}|x)\big{)}.$
By adding both sides of the above two inequalities, we get a contradiction to (10). Thus, proving that $\pi_{\theta^{*}}(y_{l}|x)<\pi_{\textrm{ref}}(y_{l}|x)$. Similarly, we can prove that $\pi_{\theta^{*}}(y_{w}|x)>\pi_{\textrm{ref}}(y_{w}|x)$, which concludes the proof. ∎
We just proved that any monotonic constraint function $\varphi$ guarantees that the learned policy, $\pi_{\theta^{*}}$, assigns a higher (lower) probability to the winner (loser) response than the one assigned to it by the reference policy $\pi_{\textrm{ref}}$. Thus, the resulting algorithm avoids the undesirable behavior of probability collapse for the winner response. It is now natural to ask what specific $\varphi$ we should use in the context of this constraint. We present two interesting candidates next.
@@ 5.1 Logarithmic Constraint $\;\varphi(x)\mathrel{\mathop{:}}=\log x$
$\log\left(\pi_{\theta}(y_{w}|x)\right)+\log\left(\pi_{\theta}(y_{l}|x)\right)=$ (12)
$\log\left(\pi_{\textrm{ref}}(y_{w}|x)\right)+\log\left(\pi_{\textrm{ref}}(y_{l}|x)\right)\ ,$
which is nice to work with empirically in light of the fact that all terms are in the log-space. Moreover, these log probabilities are already computed in DPO, which would help with efficient implementation of the corresponding C-3DPO algorithm.
Rather than using hard constraints, it is easier to compute the deviation from the constraint using either $\ell_{1}$ or $\ell_{2}$ norm, and then add it as a regularizer to the original DPO-style loss with a regularization parameter, $\lambda$, that trades-off the relative importance of the two terms. Equipping the DPO loss (5) with the logarithmic constraint (12), we obtain the following loss for C-3DPO:
$\;\;+\lambda\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{w}|x)}+\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}\right)^{2}\ .$ (13)
@@
Further, we can show that employing the logarithmic constraint (12) has a meaningful RLHF interpretation. Recall that Rafailov et al. (2023) defined
$\hat{r}_{\theta}(x,y):=\beta\log\big{(}\pi_{\theta}(y|x)/\pi_{\textrm{ref}}(y|x)\big{)},\;\forall y\in\mathcal{Y},$ (14)
as implicit reward learned during DPO training. Using this notation, we can rewrite the objective (13) simply as
@@
Under $\varphi(x)=\log x$, we solve the original RLHF problem akin to DPO, but we also incentivize that the sum of the implicit rewards for the winner and loser to stay around zero. It follows that the constraint regularizes the implicit rewards so as to avoid rewards that are (a) very large and (b) have the same sign. These two properties cannot co-exist when employing $\varphi(x)=\log x$, since doing so would yield a large magnitude inside the square and ultimately a large magnitude in the second piece of the loss. Intuitively this can hedge against probability collapse, because in the case of collapse both implicit rewards are large in magnitude and also both have negative sign, which the constraint will highly penalize.
@@ 5.2 Identity Constraint $\;\varphi(x)\mathrel{\mathop{:}}=x$
While (15) is also a plausible constraint, at first glance it is unclear how to implement it since the constraint is no longer in the log-space and is specified in terms of raw probabilities. Working with raw probabilities is prone to numerical underflow issues, thus, we would like to derive a constraint which is equivalent to (15) and operates in the log-space. To do so, we make use of the following lemma whose proof is reported in Appendix E.
@@
For any two numbers $a$ and $b$, we have
$\log(a+b)=\log a-\log\sigma(\log a-\log b)\ .$
Applying $\log$ to both sides of (15), we obtain
$\log\big{(}\pi_{\theta}(y_{w}|x)+\pi_{\theta}(y_{l}|x)\big{)}=\log\big{(}\pi_{\textrm{ref}}(y_{w}|x)+\pi_{\textrm{ref}}(y_{l}|x)\big{)}\ ,$
which can be rewritten using Lemma 5.1 as
$\log\big{(}$ $\pi_{\theta}(y_{w}|x)\big{)}-\log\sigma\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}\right)$ (16)
=\;= $\log\big{(}\pi_{\textrm{ref}}(y_{w}|x)\big{)}-\log\sigma\left(\log\frac{\pi_{\textrm{ref}}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}\right)\ .$
Moving from (15) to (16), we have rewritten the constraint entirely in the log-space, thus, avoiding numerical issues, and similar to the logarithmic constraint in Section 5.1, allowing a straightforward implementation of the corresponding C-3DPO algorithm. Equipping the DPO loss (5) with the logarithmic constraint (16), we obtain the following loss for C-3DPO:
@@
$\;\;+\lambda\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{w}|x)}+\log\sigma\left(\log\frac{\pi_{\textrm{ref}}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}\right)\right.$
$\;\;\left.-\log\sigma\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}\right)\right)^{2}\ .$ (17)
We conclude the section by noting that unlike $\varphi(x)=\log x$, deriving an RLHF interpretation under $\varphi(x)=x$ is subtle. That said, an interesting property under $\varphi(x)=x$ is that the winner probability can increase only by an amount equal to $\pi_{\textrm{ref}}(y_{l}|x)$. This means that we will not put the entire probability mass behind $y_{w}$. Thus, this constraint can serve as an additional guardrail to maintain a stochastic $\pi_{\theta}$.
@@ 6 Experiments
Figure 2: Head to head win-rate for four different implementations of C-3DPO against DPO. For each individual plot, we take the prompts from the held-out test set of Ultrafeedback Binarized and generate responses from the C-3DPO and DPO model. Then we present the prompt and two responses to Anthropic Claude Sonnet-3.5-v2 and ask which of the two answers is more helpful and honest. We then compute the win rate. Results are averaged over 10 different random seeds.
Figure 3: A comparison between C-3DPO and baselines. (Left) The two variants of C-3DPO-Log better align Zephyr-7b-SFT relative to vanilla DPO. (Right) C-3DPO variants also improve upon existing baselines that are designed to mitigate the collapse issue of DPO.
In this section, we empirically evaluate C-3DPO across 2 standard datasets and 3 initial checkpoints.
@@ 6.1 Ultrafeedback Binarized
The original Ultrafeedback dataset comprises 64k prompts, each paired with four completions from various models. To construct Ultrafeedback-Binarized, GPT-4 assigns scores based on criteria such as helpfulness. UltraFeedback_binarized was then derived by selecting the highest-scoring completion as $y_{w}$ and a randomly selected as $y_{l}$.
Figure 4: A comparison between C-3DPO and baselines with the GPT-J model on the TL;DR dataset.
Following Rasul et al. (2024), we align Zephyr-7B-SFT by performing various preference optimization algorithms on this initial checkpoint. To holistically evaluate the final model,  Rasul et al. (2024) used MT-Bench, a multi-turn benchmark that uses GPT-4 to judge models’ performance in $8$ different categories: Writing, Roleplay, Reasoning, Math, Coding, Extraction, STEM, and Humanities. In addition to MT-Bench, we also present the head to head win rate of C-3DPO against DPO by using 200 prompts from the held-out test set of Ultrafeedback-Binarized. We compute the win-rate by asking Anthropic’s Claude Sonnet-3.5-v2 which response is more helpful. The exact prompt used for Claude is in Appendix G.
Figure 5: Comparison between C-3DPO and DPO for aligning the Olmo-13B-SFT model using the Ultrafeedback-Binarized dataset.
Figure 6: Comparing C-3DPO and the baseline on the TL;DR dataset. The first four plots show the win rates of checkpoints across the first epoch against the preferred human summaries. The last figure shows the win rates of all checkpoints at different temperatures.
Recall from Section 5 that we proposed two candidates for the constraint function $\varphi$, namely the logarithmic and identity functions. Moreover, notice from (17) that we measure the deviation from the constraint using an $\ell_{2}$ penalty. Alternatively, we can measure this deviation using an $\ell_{1}$ penalty. Therefore, together, we have four specific implementations of C-3DPO, namely: C-3DPO-Log-$\ell_{1}$, C-3DPO-Log-$\ell_{2}$, $-\mathcal{I}$-$\ell_{1}$, and $-\mathcal{I}$-$\ell_{2}$. In Appendix F, we present a pseudo-code for each of these 4 variations. In Figure 2, we show the head to head win rate of each of these 4 implementations against DPO. For all 4 implementations we used the hyper-parameter $\lambda=2\times 10^{-4}$. While we did not tune $\lambda$ for each of the 4 implementations separately, doing so may increase the gap between C-3DPO and DPO.
From this result, it is clear that C-3DPO-Log is resulting in better alignment. Therefore, we use the two final checkpoints from C-3DPO-Log for MT-Bench evaluation. In Figure 3 we show the MT-Bench evaluation for C-3DPO against vanilla DPO, as well as related DPO-style algorithms that discuss the collapse of probabilities in DPO and aim to mitigate it, such as Cal-DPO (Xiao et al., 2024), SPPO (Wu et al., 2024), and DPOP (Pal et al., 2024). C-3DPO is the most competitive algorithm.
We then use C-3DPO-Log on a larger initial checkpoint, namely the Olmo-13B-SFT model from AllenAI (OLMo et al., 2024). We use the same Ultrafeedback-Binarized dataset to align this initial SFT model, and compare C-3DPO-Log against DPO. In Figure 5, we again see the superiority of C-3DPO-Log against vanilla DPO suggesting that these improvements may scale to larger models.
@@ 6.2 Reddit TL;DR
We then evaluate our proposed method on a summarization task with human-assigned scores for pairs of summaries. For this purpose, we employ the Reddit TL;DR dataset from Stiennon et al. (2022), specifically focusing on the posts under relationships and relationship_advice subreddits. We follow Amini et al. (2024) in creating the dataset and include pairs of summaries where one received a higher quality score than the other. The training and validation dataset has 572 and 102 rows respectively, each containing one Reddit post and four to twenty-four user generated summaries and their corresponding scores. During training and testing, we select the highest-scoring summary as $y_{w}$ and a randomly selected alternative as $y_{l}$.
We align the GPT-J developed by Wang and Komatsuzaki (2021) with vanilla DPO as well as C-3DPO algorithms. Specifically, we first run one SFT epoch with the Reddit TL;DR dataset on the GPT-J checkpoint, where we maximize the probability of generating the winner summary, and then take the SFT checkpoint for subsequent alignment experiments. We evaluate the final checkpoints by computing win rates against $y_{w}$ following Rafailov et al. (2023); Amini et al. (2024). We use Claude Sonnet 3.5 v2 as a judge to determine which summary is more concise and precise, and provide prompt used for Claude in the Appendix H. As shown in Figure 6, our proposed method improves upon the baseline when using higher temperatures.

### v2

@@ 5 Constrained Controlled DPO (C2-DPO)
We now present a general family of constraints to control the likelihood displacement of in-sample probabilities in DPO-style algorithms described in Section 3. These constraints can also help with the under-specified nature of these algorithms discussed in Section 4. We define the constraints on the probability mass of the winner-loser pair and use them to control how much this mass changes from the reference policy $\pi_{\textrm{ref}}$ to the target policy $\pi_{\theta}$. We then show how these constraints can be incorporated into any DPO-style loss function and propose our algorithm, which we refer to as Constrained Controlled Direct Preference Optimization (C2-DPO).
@@
The constraint, with respect to an arbitrary function $\varphi:\mathbb{R}\rightarrow\mathbb{R}$, takes the general form
$\varphi\big{(}\pi_{\theta}(y_{w}|x)\big{)}+\varphi\big{(}\pi_{\theta}(y_{l}|x)\big{)}=\varphi\big{(}\pi_{\textrm{ref}}(y_{w}|x)\big{)}+\varphi\big{(}\pi_{\textrm{ref}}(y_{l}|x)\big{)}\ .$ (9)
@@
Note that the RHS is fixed during training, so the two terms on the LHS cannot move in the same direction. We now generalize this intuition by showing that when the constraint function $\varphi$ is monotonic and added to the solution characterization of DPO-style algorithms given by
$\frac{\pi_{\theta^{*}}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{w}|x)}=\sqrt[\beta]{\frac{1-\varepsilon}{\varepsilon}}\cdot\frac{\pi_{\theta^{*}}(y_{l}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}\ ,$ (10)
then we can control the direction of the movement of probability mass for all winner-loser pairs.
@@
Proposition 5.1, whose proof is reported in Appendix D.1, shows that any monotonic constraint function $\varphi$ guarantees that the learned policy, $\pi_{\theta^{*}}$, assigns a higher (lower) probability to the winner (loser) response than the one assigned to it by the reference policy $\pi_{\textrm{ref}}$. It is natural to ask what $\varphi$ should be used in the context of this constraint, and we present two interesting candidates below.
@@ 5.1 Logarithmic Constraint $\;\varphi(x)\mathrel{\mathop{:}}=\log x$
$\log\left(\pi_{\theta}(y_{w}|x)\right)+\log\left(\pi_{\theta}(y_{l}|x)\right)=\log\left(\pi_{\textrm{ref}}(y_{w}|x)\right)+\log\left(\pi_{\textrm{ref}}(y_{l}|x)\right)\ ,$ (11)
which is nice to work with empirically in light of the fact that all terms are in the log-space. Moreover, these log probabilities are already computed in DPO, which makes the implementation of the corresponding C2-DPO algorithm more efficient.
Rather than using hard constraints, it is easier to compute the deviation from the constraint using either $\ell_{1}$ or $\ell_{2}$ norm, and then add it as a regularizer to the original DPO-style loss with a regularization parameter $\lambda$ that trades-off the relative importance of the two terms. Equipping the DPO loss (4) with the logarithmic constraint (11), we obtain the following loss for C2-DPO:
$\!\!\!\!\min_{\theta}\sum_{\mathcal{D}}\!-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{w}|x)}\!-\!\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}\right)+\lambda\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{w}|x)}\!+\!\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}\right)^{2}\!\!\!.$ (12)
@@
Further, we can show that employing the logarithmic constraint (11) has a meaningful RLHF interpretation. Recall that Rafailov et al. (2023) defined $\hat{r}_{\theta}(x,y):=\beta\log\big{(}\pi_{\theta}(y|x)/\pi_{\textrm{ref}}(y|x)\big{)},\;\forall y\in\mathcal{Y}$ as an implicit reward learned during DPO training. Using this notation, we can rewrite objective (12) simply as
@@
Under $\varphi(x)=\log x$, we solve the original RLHF problem akin to DPO, but we also incentivize the sum of the implicit rewards for the winner and loser to remain around zero. It follows that the constraint regularizes the implicit rewards so as to avoid rewards that are (a) very large and (b) have the same sign. These two properties cannot co-exist when employing $\varphi(x)=\log x$, since doing so would yield a large magnitude inside the square and ultimately a large magnitude in the second term of the loss. Intuitively, this can hedge against the likelihood displacement, because in the case of displacement both implicit rewards are large in magnitude and both have a negative sign, which the constraint will greatly penalize.
@@ 5.2 Identity Constraint $\;\varphi(x)\mathrel{\mathop{:}}=x$
While (13) is also a plausible constraint, at first glance it is unclear how to implement it since the constraint is no longer in the log-space and is specified in terms of raw probabilities. Working with raw probabilities is prone to numerical underflow issues; thus, we would like to derive a constraint which is equivalent to (13) and operates in the log-space. To do so, we make use of the following lemma whose proof is reported in Appendix D.2.
@@
For any two numbers $a$ and $b$, we have $\log(a+b)=\log a-\log\sigma(\log a-\log b)$.
Applying $\log$ to both sides of (13), we obtain $\log(\pi_{\theta}(y_{w}|x)+\pi_{\theta}(y_{l}|x))=\log(\pi_{\textrm{ref}}(y_{w}|x)+\pi_{\textrm{ref}}(y_{l}|x))$, which can be rewritten using Lemma 5.1 as
$\log\big{(}\pi_{\theta}(y_{w}|x)\big{)}-\log\sigma\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}\right)=\log\big{(}\pi_{\textrm{ref}}(y_{w}|x)\big{)}-\log\sigma\left(\log\frac{\pi_{\textrm{ref}}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}\right)\ .$ (14)
Moving from (13) to (14), we have rewritten the constraint entirely in the log-space, thus avoiding numerical issues, and similar to the logarithmic constraint in Section 5.1, allowing a straightforward implementation of the corresponding C2-DPO algorithm. Equipping the DPO loss (4) with the logarithmic constraint (14), we obtain the following loss for C2-DPO:
@@
$\qquad\qquad+\lambda\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{w}|x)}+\log\sigma\left(\log\frac{\pi_{\textrm{ref}}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}\right)-\log\sigma\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}\right)\right)^{2}\ .$ (15)
Note that unlike $\varphi(x)=\log x$, deriving an RLHF interpretation under $\varphi(x)=x$ is subtle. That said, an interesting property under $\varphi(x)=x$ is that the winner probability can increase only by an amount equal to $\pi_{\textrm{ref}}(y_{l}|x)$. This means that we will not put the entire probability mass on $y_{w}$.
@@ 6 Experiments
In this section, we present experiments on two data sets that demonstrate that C2-DPO outperforms vanilla DPO and several other baselines and delivers higher-quality final models when assessed holistically on the standard MT-Bench dataset (Zheng et al., 2023).
Figure 3: Head-to-head win-rate of C2-DPO against DPO, averaged over 10 random seeds.
Refer to captionRefer to caption
Figure 4: Left: win-rate against the preferred response in the test set of the UltraFeedback dataset. Right: Win rates on the TL;DR dataset. See Appendix for more details.
Figure 5: A comparison between C2-DPO and baselines. (Left) The two variants of C2-DPO-Log better align Zephyr-7b-SFT relative to vanilla DPO as well as the other baselines (Right).
@@ 6.1 Ultrafeedback Binarized
This dataset is comprised of 64k prompts, and a winner (preferred) and loser (rejected) continuation per prompt ($y_{w}$ and $y_{l}$). We used the standard DPO implementation published with the paper (Mitchell, 2023). Following Rasul et al. (2024), we train Zephyr-7B-SFT.
Figure 6: Left: A comparison between C2-DPO and DPO in terms of the probability of winner and loser responses during training. Right: Comparison between C2-DPO and DPO for aligning the Olmo-13B-SFT model using the Ultrafeedback-Binarized dataset.
We present the head to head win rate of C2-DPO against DPO by using 200 prompts from the held-out test set. We compute the win-rate by asking Anthropic’s Claude Sonnet-3.5-v2 which response is more helpful. The exact prompt used for Claude is in Appendix F.
Recall from Section 5 that we proposed two candidates for the constraint function $\varphi$, namely the logarithmic and identity functions. Moreover, notice from (15) that we measure the deviation from the constraint using an $\ell_{2}$ penalty. Alternatively, we can measure this deviation using an $\ell_{1}$ penalty. Altogether, we have four specific implementations of C2-DPO: C2-DPO-Log-$\ell_{1}$, C2-DPO-Log-$\ell_{2}$, $-\mathcal{I}$-$\ell_{1}$, and $-\mathcal{I}$-$\ell_{2}$. In Appendix E, we include pseudo-code for each of these 4 variations. In Figure 3, we show the head to head win rate of each of these 4 implementations against DPO, and in Figure 4 (Left), we show win-rate against the preferred response from the dataset. For all 4 implementations, we used the hyper-parameter $\lambda=2\times 10^{-4}$ and did not tune it for each of the 4 implementations separately. From this result, it is clear that all 4 variations improve upon DPO, with C2-DPO-Log being the highest performer.
Further, to holistically evaluate the final model,  Rasul et al. (2024) used MT-Bench, a multi-turn benchmark that uses GPT-4 to judge models’ performance in $8$ different categories: Writing, Roleplay, Reasoning, Math, Coding, Extraction, STEM, and Humanities. In Figure 5 we show the MT-Bench evaluation for C2-DPO against vanilla DPO, as well as related DPO-style algorithms that discuss the collapse of probabilities in DPO and aim to mitigate it, such as Cal-DPO (Xiao et al., 2024), SPPO (Wu et al., 2024), and DPOP (Pal et al., 2024). C2-DPO is the most competitive algorithm.
We then used C2-DPO-Log on a larger initial checkpoint, namely the Olmo-13B-SFT model from AllenAI OLMo et al. (2024). We compared C2-DPO-Log against DPO. In Figure 6, we see that C2-DPO-Log outperforms DPO, indicating that C2-DPO improvement may scale to larger models.
@@ 6.2 Reddit TL;DR
We then evaluate our proposed method on a summarization task with human-assigned scores for pairs of summaries. For this purpose, we employ the Reddit TL;DR dataset from Stiennon et al. (2022). We follow Amini et al. (2024) in creating the dataset and include pairs of summaries where one received a higher quality score than the other. During training and testing, we select the highest-scoring summary as $y_{w}$ and a randomly selected alternative as $y_{l}$.
We align GPT-J Wang and Komatsuzaki (2021) with vanilla DPO as well as C2-DPO algorithms. Specifically, we first run one SFT epoch with the Reddit TL;DR dataset on GPT-J, and then perform subsequent preference alignment. We evaluate the final checkpoints by computing win rates against $y_{w}$ following Rafailov et al. (2023); Amini et al. (2024). We use Claude Sonnet 3.5 v2 as a judge, and provide prompt used for Claude in the Appendix G. As shown in Figure 6, C2-DPO improves upon DPO. Moreover we show the probability of individual $y_{w}$ and $y_{l}$ in Figure 6.